# Adaptive finite elements for obstacle problems

### v1

@@ 1. Introduction
(1) $-\Delta u\geq f,\quad u\geq g,\quad(\Delta u+f)(u-g)=0,\quad u|_{\partial\Omega}=0,$
where $f$ and $g$ are given. The third condition implies that in some unknown coincidence set $\Omega_{C}\subset\Omega$ it holds $u=g$ while in $\Omega\setminus\Omega_{C}$ it holds $-\Delta u=f$. Because $\Omega_{C}$ is unknown a priori, it is challenging to construct computational meshes without relying to mesh adaptivity. In particular, more elements are usually required near the interface between the coincidence set and its complement, and the location and the shape of the interface are solution-dependent.
@@
where (.,.)(.,.)( . , . ) denotes the $L^{2}$ inner product and $H^{1}_{0}(\Omega)$ refers to the set of functions with finite energy,
@@ 2.1. Deformation and contact of a membrane
Consider a thin membrane, e.g. $D=\Omega\times[0,t]$ where $\Omega\subset\mathbb{R}^{2}$ describes the shape of the membrane on $xy$-plane and the thickness $t>0$ is very small. We wish to describe the displacement $\vec{u}:D\rightarrow\mathbb{R}^{3}$ of the membrane under a body loading $\vec{f}=(0,0,f)$. Under the kinematic assumption $\vec{u}=(0,0,u(x,y))$, infinitesimal strain $\varepsilon=\frac{1}{2}(\nabla\vec{u}+\nabla\vec{u}^{T})$, and the constitutive equation
$\sigma=2G\varepsilon+\left(K-\frac{2G}{3}\right)\mathrm{tr}\,\varepsilon I,$
the stress tensor evaluates to
$\sigma=\begin{pmatrix}0&0&Gu_{x}\\ 0&0&Gu_{y}\\ Gu_{x}&Gu_{y}&0\end{pmatrix}$
where $G>0$ is the shear modulus and $K>0$ is the bulk modulus of the material. Consequently, the governing equation is simplified:
$-\mathrm{div}\,\sigma=\vec{f}~{}\text{in $D$}\quad\Rightarrow\quad-G\Delta u=f~{}\text{in $\Omega$}.$
This is an obstacle problem if the membrane is constrained to lie above a rigid body, $u\geq g$, whose boundary is at the distance $g$ below the membrane. The corresponding constrained minimisation problem reads
$\inf_{\begin{subarray}{c}u\in H^{1}_{0}(\Omega),\\ u\geq g\end{subarray}}\frac{1}{2}(G\nabla u,\nabla u)-(f,u)$
where the membrane fixed on the boundary. The coincidence set $\Omega_{C}$ is interpreted as the zone where the rigid body and the membrane are in contact. Furthermore, the strong formulation reads: find $u:\Omega\rightarrow\mathbb{R}$ such that
(3) $-\nabla\cdot G\nabla u\geq f,\quad u\geq g,\quad(-\nabla\cdot G\nabla u+f)(u-g)=0,\quad u|_{\partial\Omega}=0,$
@@ 2.2. Elastoplastic torsion problem
As above, evaluating the infinitesimal strain and the stress tensor reveals the nonzero stress components
@@
$-\partial_{y}\sigma_{zx}+\partial_{x}\sigma_{zy}=2G\theta.$
@@
$-\Delta\phi=2G\theta~{}\text{in $\Omega$}.$
@@
(5) $-\Delta\psi\geq-2G\theta,\quad\psi\geq-\frac{\tau\delta}{\sqrt{3}},\quad(\Delta\psi-2G\theta)\left(\psi+\frac{\tau\delta}{\sqrt{3}}\right)=0,\quad\psi|_{\partial\Omega}=0.$
@@ 2.3. Hydrodynamic lubrication with cavitation
where $c_{1}$ is the bearing clearance, $e$ is the eccentricity, $10\leq e<$ 1. We wish to solve (6) on a rectangular domain $\Omega=(-R\pi,R\pi)\times(0,L)$ with periodic boundary conditions in $x=R\theta$, where $R$ is the radius of the bearing and $L$ is the width of the bearing, see Figure 2.
@@
$-\nabla\cdot\frac{d^{3}}{\mu}\nabla p\geq-6V\frac{\partial d}{\partial x},\quad p\geq p_{\text{cav}},\quad\left(\nabla\cdot\frac{d^{3}}{\mu}\nabla p-6V\frac{\partial d}{\partial x}\right)(p-p_{\text{cav}})=0,$
@@ 3. Variational formulation
$-\Delta u-\lambda=f,\quad\lambda\geq 0,\quad u\geq g,\quad\lambda(u-g)=0,\quad u|_{\partial\Omega}=0.$
@@ 4.3. Cavitation in a hydrodynamic bearing
Consequently, the experimental maximum pressure [35] is not linearly proportional to the surface velocity $V$, as one would expect based on the right hand side of (6), since the temperature is inversely proportional to the viscosity of the lubricant $\mu$. There exist separate models for thermohydrodynamics that could be utilised for studying the effect of the nonzero temperature gradient; see, e.g., Feng at al. [36]. On the other hand, we could technically use the temperature measurement data presented in Braun–Hendricks [35] to calculate a space-dependent value for the dynamic viscosity and use that value in our simulation. For simplicity, we choose a representative constant value for $\mu$ which is close to the true viscosity at the measured maximum temperature $33^{\circ}$ while the measured minimum temperature is approximately $22^{\circ}$. The cavitation pressure is a rough approximation of the minimum pressure in the pressure diagrams of [35]. We note that this problem setup can reasonably reproduce the experimental pressure at the midline, see Figure 22 for a comparison.

### v2

@@ 1. Introduction
(1) $\Delta u\geq f,\quad u\geq g,\quad(\Delta u+f)(u-g)=0,\quad u|_{\partial\Omega}=0,$
where $f$ and $g$ are given. The third condition implies that in some unknown coincidence set $\Omega_{C}\subset\Omega$ it holds $u=g$ while in $\Omega\setminus\Omega_{C}$ it holds $\Delta u=f$. Because $\Omega_{C}$ is unknown a priori, it is challenging to construct computational meshes without relying to mesh adaptivity. In particular, more elements are usually required near the interface between the coincidence set and its complement, and the location and the shape of the interface are solution-dependent.
@@
where $(.,.)$ denotes the $L^{2}$ inner product and $H^{1}_{0}(\Omega)$ refers to the set of functions with finite energy,
@@ 2.1. Deformation and contact of a membrane
Consider a flat and thin membrane, e.g. $D=\Omega\times[0,t]$ where $\Omega\subset\mathbb{R}^{2}$ describes the shape of the membrane on $xy$-plane and the thickness $t>0$ is very small. We wish to describe the out-of-plane displacement $u(x,y):\Omega\rightarrow\mathbb{R}$ of the membrane under a body loading $\vec{f}=(0,0,f)$. The governing equation is given by
$\kappa\Delta u=f~\text{in $\Omega$},$
where $\kappa$ is a uniform (pre)tension applied on the boundary of the membrane.
This becomes an obstacle problem if the membrane is constrained to lie above a rigid body, $u\geq g$, whose boundary is at the distance $g$ below the membrane. The corresponding constrained minimisation problem reads
$\inf_{\begin{subarray}{c}u\in H^{1}_{0}(\Omega),\\ u\geq g\end{subarray}}\frac{1}{2}(\kappa\nabla u,\nabla u)-(f,u)$
where the membrane is fixed on the boundary. The coincidence set $\Omega_{C}$ is interpreted as the zone where the rigid body and the membrane are in contact. Furthermore, the strong formulation reads: find $u:\Omega\rightarrow\mathbb{R}$ such that
(3) $\nabla\cdot\kappa\nabla u\geq f,\quad u\geq g,\quad(-\nabla\cdot\kappa\nabla u+f)(u-g)=0,\quad u|_{\partial\Omega}=0,$
@@ 2.2. Elastoplastic torsion problem
Evaluating the infinitesimal strain and the stress tensor reveals the nonzero stress components
@@
$\partial_{y}\sigma_{zx}+\partial_{x}\sigma_{zy}=2G\theta.$
@@
$\Delta\phi=2G\theta~\text{in $\Omega$}.$
@@
(5) $\Delta\psi\geq-2G\theta,\quad\psi\geq-\frac{\tau\delta}{\sqrt{3}},\quad(\Delta\psi-2G\theta)\left(\psi+\frac{\tau\delta}{\sqrt{3}}\right)=0,\quad\psi|_{\partial\Omega}=0.$
@@ 2.3. Hydrodynamic lubrication with cavitation
where $c_{1}$ is the bearing clearance, $e$ is the eccentricity, $0\leq e<1$. We wish to solve (6) on a rectangular domain $\Omega=(-R\pi,R\pi)\times(0,L)$ with periodic boundary conditions in $x=R\theta$, where $R$ is the radius of the bearing and $L$ is the width of the bearing, see Figure 2.
@@
$\nabla\cdot\frac{d^{3}}{\mu}\nabla p\geq-6V\frac{\partial d}{\partial x},\quad p\geq p_{\text{cav}},\quad\left(\nabla\cdot\frac{d^{3}}{\mu}\nabla p-6V\frac{\partial d}{\partial x}\right)(p-p_{\text{cav}})=0,$
@@ 3. Variational formulation
$\Delta u-\lambda=f,\quad\lambda\geq 0,\quad u\geq g,\quad\lambda(u-g)=0,\quad u|_{\partial\Omega}=0.$
@@ 4.3. Cavitation in a hydrodynamic bearing
Consequently, the experimental maximum pressure [35] is not linearly proportional to the surface velocity $V$, as one would expect based on the right hand side of (6), since the temperature is inversely proportional to the viscosity of the lubricant $\mu$. There exist separate models for thermohydrodynamics that could be utilised for studying the effect of the nonzero temperature gradient; see, e.g., Feng at al. [36]. On the other hand, we could technically use the temperature measurement data presented in Braun–Hendricks [35] to calculate a space-dependent value for the dynamic viscosity and use that value in our simulation. For simplicity, we choose a representative constant value for $\mu$ which is close to the true viscosity at the measured maximum temperature $33^{\circ}$C while the measured minimum temperature is approximately $22^{\circ}$C. The cavitation pressure is a rough approximation of the minimum pressure in the pressure diagrams of [35]. We note that this problem setup can reasonably reproduce the experimental pressure at the midline, see Figure 22 for a comparison.